OpenStudy (anonymous):

Complete the standard form of the equation -4/3y = 3 - x ?x +?y = 9

OpenStudy (ahsome):

Equation: \[\frac{-4}{3y}=3-x\]

OpenStudy (ahsome):

Right?

OpenStudy (anonymous):

Corrct

OpenStudy (ahsome):

So what do you want tot do to that equation:?

OpenStudy (anonymous):

get rid of the fraction?

OpenStudy (ahsome):

Simple. Use Cross Multiplication: \[\frac{x}{y}=\frac{a}{b}\] Is the same as: \[xb=ya\] Do you understand that?

OpenStudy (anonymous):

yea

OpenStudy (ahsome):

Equation is: \[\frac{−4}{3y}=3−x\] Is the same as: \[\frac{−4}{3y}=\frac{3−x}{1}\] Use cross multiplication: \[-4*1=3y(3-x)\] \[-4=9y-3xy\]

OpenStudy (anonymous):

??????????????????????????????????????????????????????????????????

OpenStudy (ahsome):

Whats wrong?

OpenStudy (anonymous):

The ? marks

OpenStudy (ahsome):

I know, sometimes that occur. Just refresh the page. Should work.

OpenStudy (ahsome):

Does that help?

OpenStudy (anonymous):

yes

OpenStudy (ahsome):

K. Therefore, the equation is the same as: \[−4=9y−3xy\]

OpenStudy (ahsome):

Is that it?

OpenStudy (anonymous):

?x + ?y = 9

OpenStudy (ahsome):

Hmm...

OpenStudy (anonymous):

@cj49

OpenStudy (anonymous):

@Ahsome